# Evaluating the Predictive Features of Person-Centric Knowledge Graph Embeddings: Unfolding Ablation Studies

### v1

@@ Abstract
Entity-Centric Knowledge Graphs, Graph Neural Networks, Ablation Study, Patient Representation Learning, Hospital Readmissions.
@@ 2.1 Pre-processing and Graph Neural Network Model Training
The selected task was predicting Intensive Care Unit (ICU) readmissions using the MIMIC-III dataset [9]. In this paper, we rely on a representation learning framework previously developed [5] and the Health & Social Patient-centric Ontology (HSPO) to define how PKGs are created. Readmissions are defined as any subsequent admissions within 30 days of the first ICU visit in MIMIC-III. Data was extracted such that PKGs contain information about demographic, clinical, and social context, each represented by distinct graph nodes. The demographic nodes encapsulate attributes such as age group, gender, religion, marital status, and race. For the clinical view, nodes represent diseases, received medication, and undergoing procedures. Additionally, to incorporate social aspects, information regarding employment status, household composition, and housing conditions is extracted from clinical notes using UMLS codes sampled via the MetaMap annotator [^1] MetaMap. 2016. https://www.nlm.nih.gov/research/umls/implementation_resources/metamap.html [10] and represented as nodes in the graphs.
@@ 3 Results and Discussion
Regarding the exclusion of paired clinical information, such as disease-medication or medication-procedure combinations, we observe a further impact on model efficiency, albeit not as significant as the exclusion of entire facets. This reaffirms the intercorrelation among clinical factors and underscores the need for holistic modeling approaches. Excluding clinical information completely results in significant performance degradation with up to 18.24% and 8.37% percentage decline in accuracy and F1 score (i.e., the harmonic mean of precision P and recall R: $F_{1}=2\frac{P*R}{P+R}$) respectively. Hence, this supports the fact that the clinical view of the patient may be the best predictor for the ICU readmission prediction task given the available data.

### v2

@@ Abstract
Entity-Centric Knowledge Graphs Graph Neural Networks Ablation Study Patient Representation Learning Hospital Readmissions.
@@ 2.1 Pre-processing and Graph Neural Network Model Training
The selected task was predicting Intensive Care Unit (ICU) readmissions using the MIMIC-III dataset [4]. In this paper, we rely on a representation learning framework previously developed [12] and the Health & Social Patient-centric Ontology (HSPO) to define how PKGs are created. Readmissions are defined as any subsequent admissions within 30 days of the first ICU visit in MIMIC-III. Data was extracted such that PKGs contain information about demographic, clinical, and social context, each represented by distinct graph nodes. The demographic nodes encapsulate attributes such as age group, gender, religion, marital status, and race. For the clinical view, nodes represent diseases, received medication, and undergoing procedures. Additionally, to incorporate social aspects, information regarding employment status, household composition, and housing conditions is extracted from clinical notes using UMLS codes sampled via the MetaMap annotator [^1] MetaMap. 2016.
https://www.nlm.nih.gov/research/umls/implementation_resources/metamap.html
[1] and represented as nodes in the graphs.
@@ 3 Results and Discussion
Regarding the exclusion of paired clinical information, such as medication-procedure or disease-medication combinations, we observe a further impact on model efficiency, albeit not as significant as the exclusion of entire facets. This reaffirms the intercorrelation among clinical factors and underscores the need for holistic modeling approaches. Excluding clinical information completely results in significant performance degradation with up to 18.24% and 8.37% percentage decline in accuracy and F1 score (i.e., the harmonic mean of precision P and recall R: $F_{1}=2\frac{P*R}{P+R}$) respectively. Hence, this supports the fact that the clinical view of the patient may be the best predictor for the ICU readmission prediction task given the available data.
Figure 2: Ablation study results (A, B).